You invest $5000 at 5.25% APR and is compounded quarterly. What is the investment worth after 10 years?


y=P(1+r/n)^nt


1
Expert's answer
2021-03-24T07:58:20-0400

Y= $5000


r = "\frac{4}{100}"


r/n = "\frac{5.25}{4×100}"


nt = 10 × 4 = 40


"Y=5000(1+\frac{5.25}{400})^{40}=8423.48"